Assume the Sun has been producing energy at a constant rate (3.78 x 10 ^26 J/s) over its lifetime of 4.5 billion years. How much mass has it lost creating energy over its lifetime?

____________ kg
Answer #1

E = mc2

3.78 x 10 ^26 J = m *(3*10^8m/s)^2         

m = 4.2*10^9 kg

Thus mass = 4.2*10^9kg required per second

Let us calculate for t= 4.5 billion years = 1.4*10^17s

m= 4.2 x 109 kg/s x 1.4 x 1017 s = 5.88 x 1026 kg

For 4.5 billion year mass= 5.88 x 1026 kg required